Cinnamaldehyde (M.M = 132.15 g/mol) is used as a flavoring agent. What mass of Cinnamaldehyde must be added to 175g of ethanol to give a solution whose boiling point is 82.7C. [Kb= 1.22C/m, boiling point of pure ethanol= 78.5C].
Answer #1

T = Kf*m

( 82.7 - 78.5 ) = 1.22 * m

= 3.443 = m

molality = moles / mass of solution (in kg)

= 3.443 = moles / 0.175

moles = 0.603

mass = mole * molar mass

= 0.603 * 132.15

= 79.69 grams

any query please comment

if satisfied please rate it thanks